A man weighs W on the surface of earth. What is his weight at a height equal to radius of the earth

Grade:10

1 Answers

Arun
6 years ago
Weight is inversely proportional to square of distance between man and centre of earth
W ∝ 1/d^2

Initial distance when weight = W is R
When it is at a height R distance is 2R

W₁/W₂ = (d₂/d₁)^2
W/W₂ = (2R/R)^2
W/W₂ = 4
W₂ = W/4
 
Regards
Arun (askIITians forum expert)